Q2
(a) . Given that in an economy
C = 0.8 (1-t) Y
t = 0.25
I = 900 - 50i
1
Expert's answer
2021-07-26T22:34:01-0400

The equilibrium level in the economy is derived from the IS And LM curve equilibrium position.


"Y=C+I+G"


"Y=0.8(1\u22120.25)Y+900\u221250i+800"


"Y=0.6Y+1700-50i"


"Y=0.25(1700-50i)"


"Y=4250-125i" (for IS curve equilibrium)


For LM curve equilibrium ;


"0.25Y-62.5i=500"


"Y=2000+250i"